# Tree-like process tensor contraction for automated compression of environments

### v1

@@ I Introduction
By now, several algorithms to calculate PT-MPOs have been developed. Most are aimed at solving either a generalized spin-boson model [25, 30, 31] or quantum impurities coupled to fermionic baths [32, 33]. In both cases Gaussian path integrals can be employed to derive compact expressions for the nodes of the tensor network to be contracted [34]. PT-MPOs can also be constructed for more general environments using the algorithm automated compression of environments (ACE) [28], which starts from a different tensor network: In this case, one need only assume that the environment is composed of independent modes, which could be, e.g., bosons, fermions, spins, or indeed any general non-Gaussian environment. Moreover, these modes could be time-dependently driven and themselves subject to Markovian loss or dephasing. For each mode, a PT-MPO is constructed from the respective Hamiltonian. In ACE, the individual PT-MPOs are then sequentially combined and compressed, yielding a PT-MPO encapsulating the joint influence of all modes. This algorithm is sufficiently abstract that it can be implemented by a computer code that is independent of the specific physical model at hand. Such a code can serve as a general-purpose solver to be utilized as a “black box” by the user.
@@
In this article, we demonstrate that ACE simulations can be significantly sped up by using a different contraction order of the tensor network. Concretely, we suggest to contract PT-MPOs pairwise, yielding a tree-like contraction scheme. As in Ref. [30], this scheme is enabled by preselecting degrees of freedom when combining and compressing PT-MPOs. We show several numerical examples for which this tree-like contraction scheme significantly outperforms the previous sequential contraction ACE algorithm [28]. However, we also observe that preselection results in suboptimal PT-MPO contraction, i.e. it produces PT-MPOs with larger bond dimensions than the sequential ACE algorithm. Furthermore, it accumulates larger numerical errors. We show that both of these issues can be addressed by fine-tuning convergence parameters, in particular, adjusting the number of compression sweeps per MPO combination and by choosing a scale-dependent compression threshold.
The article is structured as follows: In Section II we lay out the theory. We start with a brief summary of the ACE tensor network in Section II.1 before we discuss different schemes to contract this network in Section II.2. This includes a brief revision of the sequential contraction scheme of Ref. [28] and of the preselection approach for combining MPOs with medium to large inner bonds from Ref. [30], followed by the introduction of our tree-like contraction scheme as well as the description of strategies for fine-tuning the MPO compression. Numerical results are presented in Section III, where we discuss the model systems in Section III.1, the computation times required for simulations using different approaches in Section III.2, the corresponding numerical errors in Section III.3, and the PT-MPO compression efficacy in Section III.4. Our findings are summarized in the discussion Section IV.
@@ II.1 ACE tensor network
$\bar{\rho}_{\alpha_{n}}=$ $\sum_{\begin{subarray}{c}\alpha_{n-1},\dots,\alpha_{0}\\ \alpha^{\prime}_{n},\dots,\alpha^{\prime}_{1}\\ d_{n-1},\dots,d_{1}\end{subarray}}\prod_{l=1}^{n}\bigg{(}Q^{(\alpha_{l},\alpha^{\prime}_{l})}_{d_{l}d_{l-1}}\mathcal{M}^{\alpha^{\prime}_{l}\alpha_{l-1}}\bigg{)}\bar{\rho}_{\alpha_{0}},$ (3)
@@
Figure 1: Contraction of the ACE tensor network (a) using the sequential [(a),(b),(c),(d),(g)] and tree-like [(a),(e),(f),(g)] schemes. After combining two rows, the inner bonds are compressed by sweeps with truncated singular value decompositions. In the sequential scheme, PT-MPOs for individual environment modes are combined one by one, i.e. small contributions are added to a big PT-MPO. In the tree-like scheme, neighboring PT-MPOs are combined, so most combination steps involve the combination of PT-MPOs with small to moderate inner bond dimension. Note that panel (a) represents Eq. (2) with $e^{\mathcal{L}_{E}\Delta t}$ further decomposed as in Eq. (5). Panel (g) represents Eq. (3) with the contracted gray blocks corresponding to PT-MPO matrices $\mathcal{Q}^{(\alpha_{l},\alpha^{\prime}_{l})}_{d_{l},d_{l-1}}$.
@@
Equation (3), which is illustrated in Fig. 1(g), constitutes a series of matrix multiplications, which propagate the total density matrix from one time step to the next. However, direct application of Eq. (3) is typically infeasible because the environment Liouville space is generally extremely large. This issue can be addressed by tensor network techniques: The set of matrices $\big{\{}\mathcal{Q}^{(\alpha_{l},\alpha^{\prime}_{l})}_{d_{l},d_{l-1}}\big{\}}$ [gray comb in Fig. 1(g)] has the form of an MPO with inner bonds $d_{l}$ and outer bonds $(\alpha_{l},\alpha^{\prime}_{l})$. For such a structure, compression techniques are available, where the inner bond dimensions are reduced while not affecting the action of the total MPO on the outer bonds. This can be done by sweeping across the MPO from one end to the other while performing SVDs and eliminating degrees of freedom corresponding to small singular values [28].
@@ II.2.3 Tree-like contraction
The preselection method now allows us to explore a different contraction order, where we first combine neighboring influences of similar complexity. The resulting PT-MPOs are then further combined, leading to a tree-like overall contraction order depicted in Fig. 1 as the sequence along panels (a), (e), (f), and (g). While both, sequential ACE and the tree-like contraction, require $N_{E}-1$ combinations and compressions, the distribution of inner bond dimensions is drastically different. In particular, in the tree-like scheme, the first $N_{E}/2-1$ contractions involve combinations with the smallest possible inner bond dimensions. If neighboring modes have similar physical influences, the combined PT-MPOs should be compressible to small dimensions, which also makes the next layer of $N_{E}/4-1$ combinations and compressions efficient. Only the last few combination and compression steps involve PT-MPOs with large inner bond dimensions. This is the key argument why the tree-like contraction scheme can be expected to outperform contraction in the sequential order.
@@ III.1 Model systems
For the ACE tensor network, we discretize the frequency range $\omega=[0,7\textrm{ ps}^{-1}]$ uniformly with $N_{E}=64$ modes. The initial state of each mode is given by a thermal distribution with temperature $T=4$ K. We check convergence with respect to the Hilbert-space dimension per mode, and find $M=4$ (up to three bosonic excitations) is sufficient. A total propagation time of $t_{e}=20$ ps is discretized with steps of size $\Delta t=0.1$ ps or $\Delta t=0.01$ ps. The system Hamiltonian $H_{S}$ does not enter the PT-MPO calculation algorithm.
@@
$H_{E}^{(k)}=$ $\hbar\omega_{k}c^{\dagger}_{k}c_{k}+\hbar g_{k}(c^{\dagger}_{k}\sigma^{-}+c_{k}\sigma^{+}),$ (13)
where again $\hbar\omega_{k}$ and $g_{k}$ are mode energies and coupling constants, respectively. The operators $c^{\dagger}_{k}$ and $c_{k}$ are fermionic creation and annihilation operators, and $\sigma^{+}$ and $\sigma^{-}$ are the raising and lowering operators in a two-level system. Here, we assume that the coupling constants $g_{k}$ are identical in the (dimensionless) energy range $\omega\in[-32,32]$, which is uniformly discretized by $N_{E}=128$ modes. The magnitude of the coupling constants is chosen such that the hopping rate in the Markov limit is $\Gamma=1$. All modes are initially in the occupied state, which describes a situation where the central two-level system is energetically far below the bath’s Fermi level. A time grid spanning the time $t_{e}=5$ in steps of size $\Delta t=0.1$ is used for the calculations.
@@ III.2 Computation times
In both models, we obtain similar trends, which remain robust over the range of thresholds shown in Fig. 2: sequential ACE simulations using the original algorithm of Ref. [28] require by far the longest computation times and become impractical when high accuracies (small thresholds $\epsilon$) are needed. Using the sequential algorithm but with the preselection approach for combining rows of the tensor network already leads to massive savings in computation times. In the spin-boson example in Fig. 2(a) the computation time at $\epsilon=10^{-7}$ is reduced by a factor of about $40$, whereas we find a speed-up by a factor 6 in the fermionic model in Fig. 2(c) at $\epsilon=10^{-5}$. This is due to the fact that the larger Liouville space per mode in the spin-boson model ($M^{2}=16$ in the spin-boson model versus $M^{2}=4$ for fermions) has more potential for reduction by a prior forward sweep before combination.
Switching to the tree-like scheme, we find a further reduction of computation time with the fastest simulations being those without any fine-tuning of convergence parameters, i.e., sweeping only once ($n_{SW}=1$) per combination and using a fixed compression threshold $r=1$. The advantage over the original ACE algorithm is roughly two orders of magnitude. Compared to the method by Jørgensen and Pollock [25]—which is restricted to Gaussian environments, hence labeled as ‘Gaussian’ in Fig. 2—the more general tree-like ACE scheme requires longer computation times for those simulations with larger time steps [Fig. 2(a)]. However, when time steps are smaller [Fig. 2(b)], sequential ACE is faster than the specialized Jørgensen and Pollock algorithm. This is because the latter scales quadratically with the number of time steps, while the tree-like scheme is linear in $n$  [^4] Note however that there exist quasilinear [30] and linear-scaling [31] PT-MPO methods for Gaussian environments.
@@ III.3 Numerical error
It should be noted that very small time steps $\Delta t$ can introduce numerical artifacts. This is because in environment propagators $e^{\mathcal{L}_{E}\Delta t}\approx\mathbb{1}+\mathcal{L}_{E}\Delta t+\dots$ the information from $\mathcal{L}_{E}$ is scaled by $\Delta t$, which results in numerical values much smaller than the unity elements of the identity matrix and thus to numerical cancellation effect. Moreover, for the same overall propagation length, calculations with small times steps require many more individual combination and compression steps, and so numerical errors can accumulate. This can lead to unstable, nonconverging behavior of the tree-like scheme when error accumulation is not mitigated by a scale-dependent compression threshold [see Fig. 3(c)]. Additionally, due to the long computation times, we cannot extract reliable estimates for the compression error using sequential ACE without preselection. With preselection, the sequential ACE scheme suffers even worse numerical instabilities than the tree-like scheme, leading to unphysical results with occupations greater than unity for almost all thresholds $\epsilon$ considered [not shown in Fig. 3(c)]. We have checked that also the sequential ACE with preselection produces reliable results when an scale-dependent compression threshold $r=100$ is used. The accuracy per computation time for the fermionic bath shown in Fig. 3(d) behaves qualitative similar to that of the spin-boson model and also benefits from the use of a scale-dependent compression threshold. Interestingly, at high accuracies (corresponding to large bond dimensions) simulations using multiple sweeps $n_{\textrm{SW}}>1$ are found to be beneficial.
@@ IV Discussion
We have introduced and analyzed a novel scheme to simulate open quantum systems with general environments composed of independent modes, which is based on the algorithm ACE [28] but contracts the ACE tensor network in a different, tree-like order. The new contraction order requires a way to combine and compress PT-MPOs with large inner dimensions, for which we use the preselection approach described in Ref. [30]. Our scheme achieves about two orders of magnitude reduction in computation time compared to the original sequential ACE algorithm. The reason for this speed-up is the typical distribution of inner bond dimensions in different steps of the algorithm: most steps in the sequential scheme involve contracting an object with a small bond dimension with a PT-MPO with a large inner dimension; in contrast the tree-like scheme combines influences of similar dimensions, which are on average smaller than those of the final PT-MPO. However, the compression—as defined by the the reduction of bond dimension and numerical error—is suboptimal for the tree-like algorithm, unless the compression parameters are fine-tuned. We have explored multiple line sweeps with SVDs after combining two PT-MPOs as well as the use of a scale-dependent compression threshold; we indeed find that these strategies give a significant reduction of inner bond dimensions and of numerical errors.

### v2

@@ I Introduction
By now, several algorithms to calculate PT-MPOs have been developed. Most are aimed at solving either a generalized spin-boson model [25, 30, 31] or quantum impurities coupled to fermionic baths [32, 33]. In both cases Gaussian path integrals can be employed to derive compact expressions for the nodes of the tensor network to be contracted [34]. PT-MPOs can also be constructed for more general environments using the algorithm automated compression of environments (ACE) [28], which starts from a different tensor network: In this case, one need only assume that the environment is composed of independent modes, which could be, e.g., bosons, fermions, spins, or indeed any general non-Gaussian environment. Moreover, these modes could be time-dependently driven and themselves subject to Markovian loss or dephasing. For each mode, a PT-MPO is constructed from the respective Hamiltonian. In ACE, the individual PT-MPOs are then sequentially combined and compressed, yielding a PT-MPO encapsulating the joint influence of all modes. This algorithm is sufficiently abstract that it can be implemented by a computer code that is independent of the specific physical model at hand. Such a code can serve as a general-purpose solver to be utilized as a “black box” by the user. A C++ implementation is available and described in Ref. [35].
@@
In this article, we demonstrate that ACE simulations can be significantly sped up by using a different contraction order of the tensor network. Concretely, we suggest to contract PT-MPOs pairwise, yielding a tree-like contraction scheme. (A simple reordering of the linear sequence in which modes are contracted provides little improvement, see Appendix A.) As in Ref. [30], this scheme is enabled by preselecting degrees of freedom when combining and compressing PT-MPOs. We show several numerical examples for which this tree-like contraction scheme significantly outperforms the previous sequential contraction ACE algorithm [28]. However, we also observe that preselection results in suboptimal PT-MPO contraction, i.e. it produces PT-MPOs with larger bond dimensions than the sequential ACE algorithm. Furthermore, it accumulates larger numerical errors. We show that both of these issues can be addressed by fine-tuning convergence parameters, in particular, adjusting the number of compression sweeps per MPO combination and by choosing a scale-dependent compression threshold.
The article is structured as follows: In Section II we lay out the theory. We start with a brief summary of the ACE tensor network in Section II.1 before we discuss different schemes to contract this network in Section II.2. This includes a brief revision of the sequential contraction scheme of Ref. [28] and of the preselection approach for combining MPOs with medium to large inner bonds from Ref. [30], followed by the introduction of our tree-like contraction scheme as well as the description of strategies for fine-tuning the MPO compression. Numerical results are presented in Section III, where we discuss the model systems in Section III.1, the computation times required for simulations using different approaches in Section III.2, the corresponding numerical errors in Section III.3, and the PT-MPO compression efficacy in Section III.4. Our findings are summarized in the discussion Section IV. Further, there are three appendices: In Appendix A, we briefly investigate the impact of rearranging modes in a different sequence, while in Appendix B we provide an elementary derivation of how to obtain fermionic PT-MPOs using ACE. Finally, convergence is discussed in Appendix C using an alternative measure.
@@ II.1 ACE tensor network
$\bar{\rho}_{\alpha_{n}}=$ $\sum_{\begin{subarray}{c}\alpha_{n-1},\dots,\alpha_{0}\\ \alpha^{\prime}_{n},\dots,\alpha^{\prime}_{1}\\ d_{n-1},\dots,d_{1}\end{subarray}}\prod_{l=1}^{n}\bigg{(}\mathcal{Q}^{(\alpha_{l},\alpha^{\prime}_{l})}_{d_{l}d_{l-1}}\mathcal{M}^{\alpha^{\prime}_{l}\alpha_{l-1}}\bigg{)}\bar{\rho}_{\alpha_{0}},$ (3)
@@
Figure 1: Contraction of the ACE tensor network (a) using the sequential [(a),(b),(c),(d),(g)] and tree-like [(a),(e),(f),(g)] schemes. After combining two rows, the inner bonds are compressed by sweeps with truncated singular value decompositions. In the sequential scheme, PT-MPOs for individual environment modes are combined one by one, and so contributions with small inner bond dimensions are added to a PT-MPO with large inner bond dimension. In the tree-like scheme, neighboring PT-MPOs are combined, so most combination steps involve the combination of PT-MPOs with small to moderate inner bond dimension. Note that panel (a) represents Eq. (2) with $e^{\mathcal{L}_{E}\Delta t}$ further decomposed as in Eq. (5). Panel (g) represents Eq. (3) with the contracted gray blocks corresponding to PT-MPO matrices $\mathcal{Q}^{(\alpha_{l},\alpha^{\prime}_{l})}_{d_{l},d_{l-1}}$.
@@
Equation (3), which is illustrated in Fig. 1(g), constitutes a series of matrix multiplications, which propagate the total density matrix from one time step to the next. However, direct application of Eq. (3) is typically infeasible because the environment Liouville space is generally extremely large. This issue can be addressed by tensor network techniques: The set of matrices $\big{\{}\mathcal{Q}^{(\alpha_{l},\alpha^{\prime}_{l})}_{d_{l},d_{l-1}}\big{\}}$ [gray comb in Fig. 1(g)] has the form of an MPO with inner bonds $d_{l}$ and outer bonds $(\alpha_{l},\alpha^{\prime}_{l})$. For such a structure, compression techniques are available, where the inner bond dimensions are reduced while not affecting the action of the total MPO on the outer bonds. This can be done by sweeping across the MPO from one end to the other while performing SVDs and eliminating degrees of freedom corresponding to small singular values below a given threshold [28].
@@ II.2.3 Tree-like contraction
The preselection method now allows us to explore a different contraction order, where we first combine neighboring modes, assuming a sequence where these have similar influences (e.g., when modes are listed in order of increasing mode energies; see Appendix A for discussion of alternate orderings). The resulting PT-MPOs are then further combined, leading to a tree-like overall contraction order depicted in Fig. 1 as the sequence along panels (a), (e), (f), and (g). While both, sequential ACE and the tree-like contraction, require $N_{E}-1$ combinations and compressions, the distribution of inner bond dimensions is drastically different. In particular, in the tree-like scheme, the first $N_{E}/2-1$ contractions involve combinations with the smallest possible inner bond dimensions. If neighboring modes have similar physical influences, the combined PT-MPOs should be compressible to small dimensions, which also makes the next layer of $N_{E}/4-1$ combinations and compressions efficient. Only the last few combination and compression steps involve PT-MPOs with large inner bond dimensions. This is the key argument why the tree-like contraction scheme can be expected to outperform contraction in the sequential order.
@@ III.1 Model systems
For the ACE tensor network, we discretize the frequency range $\omega=[0,\omega_{\textrm{max}}]$ uniformly, with $\omega_{\textrm{max}}=7\textrm{ ps}^{-1}$ and using $N_{E}=64$ modes. Thus, the mode frequencies are then $\omega_{k}=(k-\frac{1}{2})\omega_{\textrm{max}}/N_{E}$ and the coupling strengths are $g_{k}=\sqrt{J(\omega_{k})\omega_{\textrm{max}}/N_{E}}$ for $k=1,\dots,N_{E}$. The initial state of each mode is given by a thermal distribution with temperature $T=4$ K. We check convergence with respect to the Hilbert-space dimension per mode, and find $M=4$ (up to three bosonic excitations) is sufficient. A total propagation time of $t_{e}=20$ ps is discretized with steps of size $\Delta t=0.1$ ps or $\Delta t=0.01$ ps. The system Hamiltonian $H_{S}$ does not enter the PT-MPO calculation algorithm.
@@
$H_{E}^{(k)}=$ $\hbar\omega_{k}c^{\dagger}_{k}c_{k}+\hbar g_{k}(c^{\dagger}_{k}c_{0}+c_{0}^{\dagger}c_{k}),$ (13)
where again $\hbar\omega_{k}$ and $g_{k}$ are mode energies and coupling constants, respectively. The operators $c^{\dagger}_{k}$ and $c_{k}$ are fermionic creation and annihilation operators, where the mode $k=0$ is the system mode while modes $k>0$ constitute the environment.
Here, we model the hopping of electrons from metallic leads to a central site. If the band width of the environment density of states is finite, the dynamics is in general non-Markovian. For concreteness, we assume that the environment density of states is centered around the system mode energy, which we shift to zero—such a shift has no effects for particle-number-conserving models. We thus discretize the (dimensionless) energy range $\omega\in[-32,32]$ uniformly with $N_{E}=128$ modes. The coupling constants $g_{k}=\sqrt{64/(2\pi N_{E})}$ are chosen such that the hopping rate in the Markov limit is $\Gamma=1$. All environment modes are initially in the occupied state, which describes a situation where the central site is energetically far below the bath’s Fermi level. A time grid spanning the time $t_{e}=5$ in steps of size $\Delta t=0.1$ is used for the calculations.
It should be noted that the derivation of ACE assumes that operators corresponding to different environment modes commute. By contrast, fermionic operators anticommute. In Appendix B, we derive how one can nevertheless obtain PT-MPOs properly accounting for the fermionic anticommutation relations using ACE. This only requires a slight modification of the input to the algorithm. This strategy can be applied irrespective of whether the contraction order is sequential or tree-like.
@@ III.2 Computation times
In both models, we obtain similar trends, which remain robust over the range of thresholds shown in Fig. 2: sequential ACE simulations using the original algorithm of Ref. [28] require by far the longest computation times and become impractical when high accuracies (small thresholds $\epsilon$) are needed. Using the sequential algorithm but with the preselection approach for combining rows of the tensor network already leads to massive savings in computation times. In the spin-boson example in Fig. 2(a) the computation time at $\epsilon=10^{-7}$ is reduced by a factor of about $40$, whereas we find a speed-up by a factor 2 in the fermionic model in Fig. 2(c) at $\epsilon=10^{-5}$. This is due to the fact that the larger Liouville space per mode in the spin-boson model ($M^{2}=16$ in the spin-boson model versus $M^{2}=4$ for fermions) has more potential for reduction by a prior forward sweep before combination.
Switching to the tree-like scheme, we find a further reduction of computation time with the fastest simulations being those without any fine-tuning of convergence parameters, i.e., sweeping only once ($n_{SW}=1$) per combination and using a fixed compression threshold $r=1$. For the spin-boson example, the advantage over the original ACE algorithm is roughly two orders of magnitude. Compared to the method by Jørgensen and Pollock [25]—which is restricted to Gaussian environments, hence labeled as ‘Gaussian’ in Fig. 2—the more general tree-like ACE scheme requires longer computation times for those simulations with larger time steps [Fig. 2(a)]. However, when time steps are smaller [Fig. 2(b)], sequential ACE is faster than the specialized Jørgensen and Pollock algorithm. This is because the latter scales quadratically with the number of time steps, while the tree-like scheme is linear in $n$  [^4] Note however that there exist quasilinear [30] and linear-scaling [31] PT-MPO methods for Gaussian environments.
@@ III.3 Numerical error
It should be noted that very small time steps $\Delta t$ can introduce numerical artifacts. This is because in environment propagators $e^{\mathcal{L}_{E}\Delta t}\approx\mathbb{1}+\mathcal{L}_{E}\Delta t+\dots$ the information from $\mathcal{L}_{E}$ is scaled by $\Delta t$, which results in numerical values much smaller than the unity elements of the identity matrix and thus to numerical cancellation effects. Moreover, for the same overall propagation length, calculations with small times steps require many more individual combination and compression steps, and so numerical errors can accumulate. This can lead to unstable, nonconverging behavior of the tree-like scheme when error accumulation is not mitigated by a scale-dependent compression threshold [see Fig. 3(c)]. Additionally, due to the long computation times, we cannot extract reliable estimates for the compression error using sequential ACE without preselection. With preselection, the sequential ACE scheme suffers even worse numerical instabilities than the tree-like scheme, leading to unphysical results with occupations greater than unity for almost all thresholds $\epsilon$ considered [not shown in Fig. 3(c)]. We have checked that also the sequential ACE with preselection produces reliable results when an scale-dependent compression threshold $r=100$ is used. The accuracy per computation time for the fermionic bath shown in Fig. 3(d) behaves qualitative similar to that of the spin-boson model and also benefits from the use of a scale-dependent compression threshold.
In Appendix C, we analyze the same situation as in Fig. 3(a) and (b) but determining the convergence error via a distance measure of MPOs. This measure is independent of the choice of system Hamiltonian and yields smoother convergence curves, but numerical cancellation prevents its use very close to convergence.
@@ IV Discussion
We have introduced and analyzed a novel process tensor contraction scheme, with which the ACE algorithm [28] for the simulation of open quantum systems can be considerably improved. The new, tree-like contraction order requires a way to combine and compress PT-MPOs with large inner dimensions, for which we use the preselection approach described in Ref. [30]. Our scheme achieves about two orders of magnitude reduction in computation time compared to the original sequential ACE algorithm. The reason for this speed-up is the typical distribution of inner bond dimensions in different steps of the algorithm: most steps in the sequential scheme involve contracting an object with a small bond dimension with a PT-MPO with a large inner dimension; in contrast the tree-like scheme combines influences of similar dimensions, which are on average smaller than those of the final PT-MPO. However, the compression—as defined by the the reduction of bond dimension and numerical error—is suboptimal for the tree-like algorithm, unless the compression parameters are fine-tuned. We have explored multiple line sweeps with SVDs after combining two PT-MPOs as well as the use of a scale-dependent compression threshold; we indeed find that these strategies give a significant reduction of inner bond dimensions and of numerical errors.